# Neural Exploratory Landscape Analysis

### v2

@@ Abstract
Recent research in Meta-Black-Box Optimization (MetaBBO) have shown that meta-trained neural networks can effectively guide the design of black-box optimizers, significantly reducing the need for expert tuning and delivering robust performance across complex problem distributions. Despite their success, a paradox remains: MetaBBO still rely on human-crafted Exploratory Landscape Analysis features to inform the meta-level agent about the low-level optimization progress. To address the gap, this paper proposes Neural Exploratory Landscape Analysis (NeurELA), a novel framework that dynamically profiles landscape features through a two-stage, attention-based neural network, executed in an entirely end-to-end fashion. NeurELA is pre-trained over a variety of MetaBBO algorithms using a multi-task neuroevolution strategy. Extensive experiments show that NeurELA achieves consistently superior performance when integrated into different and even unseen MetaBBO tasks and can be efficiently fine-tuned for further performance boost. This advancement marks a pivotal step in making MetaBBO algorithms more autonomous and broadly applicable. The source code of NeurELA can be accessed at https://anonymous.4open.science/r/Neur-ELA-303C.
@@ 1 Introduction
Black-Box Optimization (BBO) is a discipline in mathematical optimization characterized by the absence of both the closed-form expression of the target objective function and its derivative during the optimization process. Traditional BBO methods, such as Genetic Algorithm [1], Evolutionary Strategy [2], and Bayesian Optimization [3], have been extensively studied in the past half century. They are applied to address a broad range of black-box scenarios, including automated machine learning [4], bioinformatics [5], prompt tuning in large language models [6]. However, these methods often require manual tuning with deep expertise to ensure the optimal performance on unseen tasks. Recent advances in Meta-Black-Box Optimization (MetaBBO) [7, 8, 9, 10, 11] leverage meta-learning [12] to streamline the traditional labour-intensive tuning process across different problem distributions. This effort not only reduces the need for expert-level knowledge, but also enhances generalization towards unseen problems. Generally, MetaBBO employs a bi-level optimization paradigm, wherein the meta-level control policy (often implemented as a neural network) is optimized according to the feedback from the low-level optimization process [13], as depicted on the left side of Figure 1.
@@
Existing MetaBBO algorithms widely adopt Exploratory Landscape Analysis (ELA) [14, 15] to profile the low-level optimization status. While the approaches provide a comprehensive understanding of problem and algorithm characteristics, they still face several limitations. 1) Many features derived from ELA are highly correlated with each other [16] and exhibit strong sensitivity to random sampling strategies [17]. This necessitates an additional step of feature selection for the given MetaBBO scenario, requiring a certain level of expert knowledge. 2) The landscape features given by ELA are more like an “offline” portray of the target optimization problem. This is somewhat incompatible with the dynamic nature of MetaBBO paradigm, which requires both the “offline” problem properties and the “online” optimization progress. 3) Computing some of these features, particularly those related to convexity and local optimum properties, consume additional function evaluations (FEs). The consumption of the computational resource, initially reserved for the low-level optimization, ultimately degrades the final performance of the MetaBBO algorithms.
@@
2) We instantiate the landscape analyser $\Lambda$ through a neural network model parameterized by $\theta$. Specifically, $\Lambda_{\theta}$ is implemented using a two-stage attention-based neural network. Our design enables flexible integration with most existing MetaBBO algorithms as an effective replacement for their traditional hand-crafted landscape analysis mechanisms. Given the non-differential nature of our objective in BBO, we adopt an Evolution Strategy [18, 19] to train $\Lambda_{\theta}$. This training is carried out within a multitask neuroevolution [20] paradigm, aiming to optimize $\theta$ to maximize the expected performance across representative MetaBBO tasks.
@@
Through extensive experiments, we validate the effectiveness of the proposed NeurELA framework, as illustrated on the right side of Figure 1. These results demonstrate that it is possible to pre-train an universal neural landscape analyser $\Lambda_{\theta}$ over a group of MetaBBO tasks and then generalize it to unseen ones. Our work pioneers the integration of automatic feature extraction in MetaBBO, transforming these methods into fully end-to-end systems. This advancement significantly streamlines optimization processes, enhancing efficiency and adaptability across related fields, and reducing the need for manual expertise.
@@ 2 Related Works
Meta-Black-Box Optimization. Existing MetaBBO paradigms can be classified into following categories with respect to role they play at the low-level optimization process: 1) Hyper-parameters configurator: this line of works meta-trained neural network-based policy to dynamically configure the hyper-parameters of the backend BBO optimizer, e.g., the scale factor ($F$) and crossover rate ($Cr$) in Differential Evolution (DE) algorithm [21, 22]. Initial researches (e.g., RLPSO [23]) employed simple MLP structures to parameterize the meta-level policy, which was enhanced by the latest works such as GLEET [24], which facilitated Transformer [25] architecture to boost the sequential decision making along the low-level optimization. 2) Update rule selector: works in this line flexibly select desired update rules for the low-level optimization process. Several initial works learns to select one mutation operator from a pre-defined DE mutation operator pool to attain diverse optimization behaviours [26, 27, 22]. Then the research scope extends to the refinement of the operators pool for optimizing more complex problems [28], and dynamically selecting the entire algorithm[10]. 3) Complete optimizer generator: works in this line aims to meta-learning the neural network-based policy as a complete optimizer. Lange et al. [8, 11] proposed meta-training several attention blocks as the complete updating process of CMAES [29]  , which was followed by Song et al. [30], proposing meta-training Transformer-based language model as an entire BBO method. Such end-to-end deep learning systems encountered interpretability issue. To this end, SYMBOL [9] proposed generating symbolic update rules for the low-level optimization process, promoting the interpretable and automatic discovery of black-box method. GLHF [31] carefully design a Transformer-based neural network that imitates the evolutionary operators in DE algorithm. As for the designs of optimization status extraction, a great majority of existing MetaBBO algorithms borrow the idea from well-known landscape analysis researches [14, 15]. To make those landscape features compatible with the MetaBBO paradigm, existing MetaBBO algorithms more or less facilitated feature selection and feature aggregation procedure. Take a recent work GLEET [24] as an example, it aggregates the representative ELA features [15] into a $9$-dimensional high-level profiling features, which helps reduce the training complexity and enhance the overall optimization performance. However, we have to note that such hand-crafted feature processing requires deep expertise in optimization.
@@
Landscape Analysis. Landscape Analysis originates from the emergence of Automated Algorithm Selection (AAS) [32, 33], which selects an optimal algorithm out of a collection of algorithms for any given optimization problem. Generally, Landscape analysis is formulated as a function: $\Lambda:(X,y)\rightarrow s$, that maps the sampled candidates $X$ and their corresponding objective values $y$ to a feature vector $s$, each dimension of which indicates some problem-specific properties of the target optimization problem. The feature vector $s$ can be subsequently fed into machine learning algorithms (e.g., Support Vector Machines [34] and Random Forests [35]) to address some specific algorithmic tasks [36]. Over the years, a wide range of Landscape Analysis feature groups have been developed, and structurally summarized in [14, 37, 38, 39], of which the mostly utilized ones are: 1) Fitness Distance Correlation [40] includes six features profiling the distances of the sampled candidates both in decision space and the objective space, as well as the correlation in-between the two spaces. 2) Dispersion [41] features describe the degree of dispersity for the best quantiles of the sampled candidates. 3) Classical ELA [15] proposed using six groups of low-level metrics such as local search, skewness of objective space and approximated curvature in (second) first-order as a comprehensive abstract of basic landscape properties (e.g., global structure, multimodality, separability). Generally, classic ELA requires adequate samples and enormous computational efforts to conclude all of its fifty low-level feature values. 4) Information Content [42] comprises five features about the smoothness, ruggedness and neutrality of an optimization problem. It first applies random walk sampling (e.g., Metropolis Random Walk [43]) for each of the sampled candidates, then calculates its features by statistic metrics. 5) Nearest Better Clustering [44] is a collection of features that describe the distributional correlation of the nearest better neighbors and the nearest neighbors of the sampled candidates.
@@
While these hand-crafted landscape analysis features provide a systematic understanding of optimization problems, they inevitably hold high computing complexity hence consume additional computational resources. Besides, the intricate correlations in-between these feature groups [16] and their sensitivities to the random sampling strategy [17] necessitate the expert-level feature selection and feature aggregation procedures. To address these issues, several early-stage studies have recently proposed learning neural networks (e.g., convolutional neural network [45] and point cloud transformer [46]) as the feature mapping function. [47] first proposed using Deep-CNN to extract landscape features from fitness map of moderate initial sample points. The neural network is trained to correctly predict algorithm configuration of modular CMA-ES algorithms. However, solely feeding the fitness map might introduce information loss. To address such issue, [48] propose a novel point cloud transformer (PCT) based approach, which embeds sample points information and the fitness map together as the input. This work generally achieves superior landscape feature prediction accuracy. Another technical bottleneck of [47] is the supervised training, which might leads the learned landscape representation overfit to a certain AAS/AAC task. To refine this issue, [49] propose self-supervised learning through AutoEncoder, which adopted fitness map as the target to be reconstructed. More recently, [50] propose DeepELA to adrress both the infomation loss and the supervised trainng issues. It introduces MHA and trains it by a contrastive learning loss InfoNCE to ensure aligned representation across different sampling strategies. In this paper, our NeurELA presents several novelty compared with the above learning-based ELA frameworks: a) NeurELA aims for MetaBBO tasks yet the previous works generally aim for classical AS or AC tasks. MetaBBO tasks are more akin to Dynamic AS or AC tasks (DAC, DAS), where the configuration decision have to be customized in each generation of optimization of a given problem. b) Unlike previous works, which provide static and invariant problem properties, NeurELA offers dynamic optimization status information. c) compared with [50], our architecture designs do not rely on a pre-defined dimension upper bound, making it more generic and versatile for various MetaBBO tasks.
@@ 3.1 Framework
In accordance with the MetaBBO protocol, each algorithm $\mathbb{A}_{k}$ is composed of three fundamental elements: the meta-level neural policy $\Pi_{\phi}$, the low-level optimizer $\Gamma$, and the initial landscape feature analyser, denoted as $\Lambda_{0}$. As illustrated in Figure 1, the meta-training of $\mathbb{A}_{k}$ adheres to a bi-level optimization paradigm. Given a problem instance $f$ sampled from $\mathbb{D}_{k}$, at each optimization step $t$, the low-level BBO method $\Gamma$ interacts with $f$ to attain a raw observation $\{X^{t}_{i},y^{t}_{i}\}_{i=1}^{m}$ of the current population of candidates solutions, denoted as $o^{t}$, where $m$ is the population size in population-based search methods [51, 52, 1]. Subsequently, the observation $o^{t}$ is further processed by the original landscape analyser, $\Lambda_{0}$, to extract optimization status features $s^{t}=\Lambda_{0}\left(o^{t}\right)$. Based on these features, the meta-level neural policy $\Pi_{\phi}$ dictates a design choice $c^{t}$ from the configuration space $\mathbb{C}$ conditioned on the current optimization status: $c^{t}=\Pi_{\phi}(s^{t})$. Then, the low-level BBO method $\Gamma$ loads $c^{t}$ to optimize $f$ for next time step $t+1$. The meta-performance $r^{t}$ is calculated by evaluating the performance gain achieved by implementing $c_{t}$. The ultimate goal of the meta-training phase is to maximize the overall meta-objective $F_{\text{metaBBO}}$, defined as:
@@
where $P=\left|\mathbb{D}_{\text{test}}\right|$ is the number of test problems (the dataset $\mathbb{D}_{k}$ is divided into two distinct subsets: $\mathbb{D}_{\text{train}}$ for training and $\mathbb{D}_{\text{test}}$ for testing), $Q=50$ represents the number of independent runs to obtain average performance, $f_{p,q}^{*}$ denotes the objective value achieved by testing $\mathbb{A}_{k}^{\prime}$ on the $p$-th problem instance and $q$-th run, and $\mu_{p}$ and $\sigma_{p}$ are the mean and standard deviation of the objective values obtained by testing $\mathbb{A}_{k}$ on the $p$-th problem over $50$ independent runs. Note that we follow the Z-score normalization design in [13] to evaluate the performance gain $Z_{p,q}$.

### v3

@@ Abstract
Recent research in Meta-Black-Box Optimization (MetaBBO) have shown that meta-trained neural networks can effectively guide the design of black-box optimizers, significantly reducing the need for expert tuning and delivering robust performance across complex problem distributions. Despite their success, a paradox remains: MetaBBO still rely on human-crafted Exploratory Landscape Analysis features to inform the meta-level agent about the low-level optimization progress. To address the gap, this paper proposes Neural Exploratory Landscape Analysis (NeurELA), a novel framework that dynamically profiles landscape features through a two-stage, attention-based neural network, executed in an entirely end-to-end fashion. NeurELA is pre-trained over a variety of MetaBBO algorithms using a multi-task neuroevolution strategy. Extensive experiments show that NeurELA achieves consistently superior performance when integrated into different and even unseen MetaBBO tasks and can be efficiently fine-tuned for further performance boost. This advancement marks a pivotal step in making MetaBBO algorithms more autonomous and broadly applicable. The source code of NeurELA can be accessed at https://github.com/GMC-DRL/Neur-ELA.
@@ 1 Introduction
Black-Box Optimization (BBO) is a discipline in mathematical optimization characterized by the absence of both the closed-form expression of the target objective function and its derivative during the optimization process. Traditional BBO methods, such as Genetic Algorithm (Holland, 1975), Evolutionary Strategy (Hansen & Ostermeier, 2001a), and Bayesian Optimization (Shahriari et al., 2015), have been extensively studied in the past half century. They are applied to address a broad range of black-box scenarios, including automated machine learning (Akiba et al., 2019), bioinformatics (Tsaban et al., 2022), prompt tuning in large language models (Guo et al., 2024b). However, these methods often require manual tuning with deep expertise to ensure the optimal performance on unseen tasks. Recent advances in Meta-Black-Box Optimization (MetaBBO) (Xue et al., 2022; Lange et al., 2023; Chen et al., 2024; Guo et al., 2024a; Lange et al., 2024) leverage meta-learning (Finn et al., 2017) to streamline the traditional labour-intensive tuning process across different problem distributions. This effort not only reduces the need for expert-level knowledge, but also enhances generalization towards unseen problems. As depicted on the left side of Figure 1, MetaBBO generally employs a bi-level optimization paradigm (Ma et al., 2023). In the meta level, a control policy (often implemented as a neural network) is maintained to dictate timely algorithm configuration for the low-level BBO optimizer according to the low-level optimization status. Once obtained the algorithm configuration, the low-level BBO optimizer optimizes the target BBO problem by that configuration and returns a feedback to the meta-level policy the performance gain feedback. The meta-objective of MetaBBO is to meta-learn an optimal control policy which could achieve maximum accumulated performance gain across a problem distribution.
@@
Existing MetaBBO algorithms widely adopt Exploratory Landscape Analysis (ELA) (Wright et al., 1932; Mersmann et al., 2011) to profile the low-level optimization status. While the approaches provide a comprehensive understanding of problem and algorithm characteristics, they still face several limitations. 1) Many features derived from ELA are highly correlated with each other (Renau et al., 2019) and exhibit strong sensitivity to random sampling strategies (Renau et al., 2020). This necessitates an additional step of feature selection for the given MetaBBO scenario, requiring a certain level of expert knowledge. 2) The landscape features given by ELA are more like an “offline” portray of the target optimization problem. This is somewhat incompatible with the dynamic nature of MetaBBO paradigm, which requires both the “offline” problem properties and the “online” optimization progress. 3) Computing some of these features, particularly those related to convexity and local optimum properties, consume additional function evaluations (FEs). The consumption of the computational resource, initially reserved for the low-level optimization, ultimately degrades the final performance of the MetaBBO algorithms.
@@
2) We instantiate the landscape analyser $\Lambda$ through a neural network model parameterized by $\theta$. Specifically, $\Lambda_{\theta}$ is implemented using a two-stage attention-based neural network. Our design enables flexible integration with most existing MetaBBO algorithms as an effective replacement for their traditional hand-crafted landscape analysis mechanisms. Given the non-differential nature of our objective in BBO, we adopt an Evolution Strategy (Hansen & Ostermeier, 2001b; Li et al., 2018) to train $\Lambda_{\theta}$. This training is carried out within a multitask neuroevolution (Miikkulainen et al., 2024) paradigm, aiming to optimize $\theta$ to maximize the expected performance across MetaBBO tasks.
@@
Through extensive experiments, we validate the effectiveness of NeurELA, as illustrated on the right side of Figure 1. These results demonstrate that it is possible to pre-train a universal neural landscape analyser $\Lambda_{\theta}$ over a group of MetaBBO tasks and then generalize it to unseen ones. Our work pioneers the integration of automatic feature extraction in MetaBBO, transforming these methods into fully end-to-end systems. This advancement significantly streamlines optimization processes, enhancing efficiency and adaptability across related fields, and reducing the need for manual expertise.
@@ 2 Related Works
Meta-Black-Box Optimization. Existing MetaBBO paradigms can be classified into following categories with respect to role they play at the low-level optimization process: 1) Hyper-parameters configurator: this line of works meta-trained neural network-based policy to dynamically configure the hyper-parameters of the backend BBO optimizer, e.g., the scale factor ($F$) and crossover rate ($Cr$) in Differential Evolution (DE) algorithm (Sun et al., 2021; Tan et al., 2022). Initial researches (e.g., RLPSO (Wu & Wang, 2022)) employed simple MLP structures to parameterize the meta-level policy, which was enhanced by the latest works such as GLEET (Ma et al., 2024), which facilitated Transformer (Vaswani et al., 2017) architecture to boost the sequential decision making along the low-level optimization. 2) Update rule selector: works in this line flexibly select desired update rules for the low-level optimization process. Several initial works learns to select one mutation operator from a pre-defined DE mutation operator pool to attain diverse optimization behaviours (Sharma et al., 2019; Tan & Li, 2021; Tan et al., 2022). Then the research scope extends to the refinement of the operators pool for optimizing more complex problems (Lian et al., 2024), and dynamically selecting the entire algorithm(Guo et al., 2024a). 3) Complete optimizer generator: works in this line aims to meta-learning the neural network-based policy as a complete optimizer. Lange et al. (Lange et al., 2023; 2024) proposed meta-training several attention blocks as the complete updating process of CMAES (Hansen et al., 2003)  , which was followed by Song et al. (Song et al., 2024), proposing meta-training Transformer-based language model as an entire BBO method. Such end-to-end deep learning systems encountered interpretability issue. To this end, SYMBOL (Chen et al., 2024) proposed generating symbolic update rules for the low-level optimization process, promoting the interpretable and automatic discovery of black-box method. GLHF (Li et al., 2024) carefully design a Transformer-based neural network that imitates the evolutionary operators in DE algorithm. As for the designs of optimization status extraction, a great majority of existing MetaBBO algorithms borrow the idea from well-known landscape analysis researches (Wright et al., 1932; Mersmann et al., 2011). To make those landscape features compatible with the MetaBBO paradigm, existing MetaBBO algorithms more or less facilitated feature selection and feature aggregation procedure. Take a recent work GLEET (Ma et al., 2024) as an example, it aggregates the representative ELA features (Mersmann et al., 2011) into a $9$-dimensional high-level profiling features, which helps reduce the training complexity and enhance the overall optimization performance. However, we have to note that such hand-crafted feature processing requires deep expertise in optimization.
@@
Landscape Analysis. Landscape Analysis originates from the emergence of Automated Algorithm Selection (AAS) (Rice, 1976; Kerschke et al., 2019), which selects an optimal algorithm out of a collection of algorithms for any given optimization problem. Generally, Landscape analysis is formulated as a function: $\Lambda:(X,y)\rightarrow s$, that maps the sampled candidates $X$ and their corresponding objective values $y$ to a feature vector $s$, each dimension of which indicates some problem-specific properties of the target optimization problem. The feature vector $s$ can be subsequently fed into machine learning algorithms (e.g., Support Vector Machines (Cortes & Vapnik, 1995) and Random Forests (Breiman, 2001)) to address some specific algorithmic tasks (Kerschke & Trautmann, 2019a). Over the years, a wide range of Landscape Analysis feature groups have been developed, and structurally summarized in (Wright et al., 1932; Pitzer & Affenzeller, 2012; Malan & Engelbrecht, 2013; Kerschke & Trautmann, 2019b), of which the mostly utilized ones are: 1) Fitness Distance Correlation (Jones et al., 1995) includes six features profiling the distances of the sampled candidates both in decision space and the objective space, as well as the correlation in-between the two spaces. 2) Dispersion (Lunacek & Whitley, 2006) features describe the degree of dispersity for the best quantiles of the sampled candidates. 3) Classical ELA (Mersmann et al., 2011) proposed using six groups of low-level metrics such as local search, skewness of objective space and approximated curvature in (second) first-order as a comprehensive abstract of basic landscape properties (e.g., global structure, multimodality, separability). Generally, classic ELA requires adequate samples and enormous computational efforts to conclude all of its fifty low-level feature values. 4) Information Content (Muñoz et al., 2014) comprises five features about the smoothness, ruggedness and neutrality of an optimization problem. It first applies random walk sampling (e.g., Metropolis Random Walk (Hastings, 1970)) for each of the sampled candidates, then calculates its features by statistic metrics. 5) Nearest Better Clustering (Kerschke et al., 2015) is a collection of features that describe the distribution of the nearest better neighbors and the nearest neighbors of the sampled candidates.
@@
While these hand-crafted landscape analysis features provide a systematic understanding of optimization problems, they inevitably hold high computing complexity hence consume additional computational resources. Besides, the intricate correlations in-between these feature groups (Renau et al., 2019) and their sensitivities to the random sampling strategy (Renau et al., 2020) necessitate the expert-level feature selection and feature aggregation procedures. To address these issues, several early-stage studies have recently proposed learning neural networks (e.g., convolutional neural network (Prager et al., 2021a) and point cloud transformer (Seiler et al., 2022a)) as the feature mapping function. Prager et al. (2021b) first proposed using Deep-CNN to extract landscape features from fitness map of moderate initial sample points. The neural network is trained to correctly predict algorithm configuration of modular CMA-ES algorithms. However, solely feeding the fitness map might introduce information loss. To address such issue, Seiler et al. (2022b) propose a novel point cloud transformer (PCT) based approach, which embeds sample points information and the fitness map together as the input. This work generally achieves superior landscape feature prediction accuracy. Another technical bottleneck of (Prager et al., 2021b) is the supervised training, which might leads the learned landscape representation overfit to a certain AAS/AAC task. To refine this issue, van Stein et al. (2023) propose self-supervised learning through AutoEncoder, which adopted fitness map as the target to be reconstructed. More recently, Seiler et al. (2024) propose DeepELA to adrress both the infomation loss and the supervised trainng issues. It introduces MHA and trains it by a contrastive learning loss (InfoNCE) to ensure aligned representation across different sampling strategies. In this paper, our NeurELA presents several novel contributions compared with the above learning-based ELA frameworks: a) NeurELA is designed for dynamic MetaBBO tasks, enabling dynamic algorithm configuration and selection, whereas previous studies focus on static profiling of a given problem. b) NeurELA introduces a scalable embedding strategy that removes dimensional constraints in DeepELA and employs a two-stage attention mechanism for extracting features across sample points and problem dimensions, making it more generic and versatile for various MetaBBO tasks. For more detailed differences, please refer to Appendix B.2.
@@ 3.1 Framework
In accordance with the MetaBBO protocol, each algorithm $\mathbb{A}_{k}$ is composed of three fundamental elements: the meta-level neural policy $\Pi_{\phi}$, the low-level optimizer $\Gamma$, and the initial landscape feature analyser, denoted as $\Lambda_{0}$. As illustrated in Figure 1, the meta-training of $\mathbb{A}_{k}$ adheres to a bi-level optimization paradigm. Given a problem instance $f$ sampled from $\mathbb{D}_{k}$, at each optimization step $t$, the low-level BBO method $\Gamma$ interacts with $f$ to attain a raw observation $\{X^{t}_{i},y^{t}_{i}\}_{i=1}^{m}$ of the current population of candidates solutions, denoted as $o^{t}$, where $m$ is the population size in population-based search methods Storn & Price (1997); Kennedy & Eberhart (1995); Holland (1975). Subsequently, the observation $o^{t}$ is further processed by the original landscape analyser, $\Lambda_{0}$, to extract optimization status features $s^{t}=\Lambda_{0}\left(o^{t}\right)$. Based on these features, the meta-level neural policy $\Pi_{\phi}$ dictates a design choice $c^{t}$ from the configuration space $\mathbb{C}$ conditioned on the current optimization status: $c^{t}=\Pi_{\phi}(s^{t})$. Then, the low-level BBO method $\Gamma$ loads $c^{t}$ to optimize $f$ for next time step $t+1$. The meta-performance $r^{t}$ is calculated by evaluating the performance gain achieved by implementing $c_{t}$. The ultimate goal of the meta-training phase is to maximize the overall meta-objective $F_{\text{metaBBO}}$, defined as:
@@
where $P=\left|\mathbb{D}_{\text{test}}\right|$ is the number of test problems (the dataset $\mathbb{D}_{k}$ is divided into two distinct subsets: $\mathbb{D}_{\text{train}}$ for training and $\mathbb{D}_{\text{test}}$ for testing), $Q=50$ represents the number of independent runs to obtain average performance, $f_{p,q}^{*}$ denotes the objective value achieved by testing $\mathbb{A}_{k}^{\prime}$ on the $p$-th problem instance and $q$-th run, and $\mu_{p}$ and $\sigma_{p}$ are the mean and standard deviation of the objective values obtained by testing $\mathbb{A}_{k}$ on the $p$-th problem over $50$ independent runs. Note that we follow the Z-score normalization design in Ma et al. (2023) to evaluate the performance gain $Z_{p,q}$.